OpenStudy (anonymous):

7x - 3y = 4 2x - 4y = 1 Which of the following systems is not the same as solving the system shown? 28x - 12y = 16 and -6x + 12y = -3 14x - 6y = 4 and -14x + 28y = 1 -28x + 12y = -16 and 28x - 56y = 14

OpenStudy (anonymous):

7x−3y=4;2x−4y=1 Step: Solve 7x−3y=4 for x: 7x−3y+3y=4+3y(Add 3y to both sides) 7x=3y+4 7x7=3y+47(Divide both sides by 7) x=37y+47 Step: Substitute (37y+47) for x in 2x−4y=1: 2x−4y=1 2(37y+47)−4y=1 −227y+87=1(Simplify both sides of the equation) −227y+87+−87=1+−87(Add (-8)/7 to both sides) −227y=−17 −227y−227=−17−227(Divide both sides by (-22)/7) y=122 Step: Substitute (122) for y in x=37y+47: x=37y+47 x=37(122)+47 x=1322(Simplify both sides of the equation) Answer: x=1322 and y=122

OpenStudy (welshfella):

you can do it this way look at each equation in one of the options if the equations are the result of multiplying each term by the same constant then the equations are the same do the above with each of the 3 sets

OpenStudy (welshfella):

for example try option 1 28x - 12y = 16 is obtained by multiplying 7x -3y = 4 by 4

OpenStudy (welshfella):

and -6x + 12y = -3 is obtained by multiplying 2x - 4y = 1 by -3 so the first option is the same as the original equations

OpenStudy (welshfella):

so try options 2 and 3 the same way

OpenStudy (welshfella):

to start you off the first equation in second set 7x - 3y = 4 is multiplied by 2 - did they do it correctly?

OpenStudy (welshfella):

oh!! - talking to myself again!